OpenStudy (anonymous):

Solve: −4(x + 1) − 3 = −3(x − 4)

OpenStudy (eric_d):

expand

OpenStudy (eric_d):

-4x-4-3=-3x+12

OpenStudy (eric_d):

then

OpenStudy (eric_d):

rearrange

OpenStudy (eric_d):

x to LHS

OpenStudy (eric_d):

LHS= LEFT HAND SIDE

OpenStudy (anonymous):

you will need to distribute

OpenStudy (anonymous):

−4(x + 1) − 3 = −3(x − 4) -4x - 4 -3 = -3x + 12 -4x - 7 = -3x + 12 -x = 19 x = -19

OpenStudy (anonymous):

thank you guys! (: @FriedRice @eric_d

OpenStudy (eric_d):

yw